OpenStudy (anonymous):

Consider 2x<100 a. Determine an expression for the cost. b. State the domain of this function. c. Give the full name of this kind of function

OpenStudy (anonymous):

Thanks in advance guys!

OpenStudy (anonymous):

Sorry, I wasn't exactly on the question.\[C(x) = \left\{\begin{array}{r} 2x, x\le 100 \\2*100 + 1.75x, x>100\end{array}\right. \]

OpenStudy (anonymous):

there are two cases, so you have a combination of equations for it one casse is for the first 100 chips, the second case is for more than 100

OpenStudy (anonymous):

and Ish is spot on, but you need to rewrite that function using p instead of x